# A hybrid numerical method for elastic wave propagation in discontinuous media with complex geometry

### v1

@@ 3.2 Stability of the coupled method
It is proven in A that the conditions above are met for the choice $\chi_{2}=2\mu_{u}+\lambda_{u}$ if the penalty parameter $\beta_{v}$ satisfies the bound $\beta_{u}>\frac{1}{4C^{\prime}}$, as desired. Hence we have proven that (24) is non-negative and therefore we can conclude that $U_{2}\geq 0$ under the assumption $\beta_{u}\leq\frac{1}{4C^{\prime}}$ which together with (22) being non-negative if $\alpha_{v}\geq\frac{1}{4C}$ proves that
$\frac{1}{2}\frac{d}{dt}(\rho_{u}\boldsymbol{u}_{t}^{T}\mathcal{H}\boldsymbol{u}_{t}+(\rho_{v}\boldsymbol{v}_{t},\boldsymbol{v}_{t})_{\Omega_{v}})=0,$
@@
It should be noted that the proof will fail at (16) if we remove the assumption of full compatibility as we cannot obtain a total time derivative. This assumption can be relaxed for the wave equation [1]. However, we see that the assumption of norm-compatible projection pairs from [3] is fundamental to be able to derive the estimate.
@@ 4 Numerical results
In this section, we present numerical examples to demonstrate accuracy and stability properties of the FD-DG scheme developed for the elastic wave equation. [^1] The MATLAB codes used to perform the experiments are available at https://github.com/AndreasGranath/Elastic-Waves-FD-DG. First, we examine how the selection of free parameters influences the scheme’s stability and time-step restriction. Then, we conduct a series of benchmarking examples to demonstrate the stability and convergence properties of the developed method. This is demonstrated in three scenarios: first, utilizing a manufactured solution in a homogeneous medium with straight boundary and interface; then, introducing a curved boundary within the same medium; and lastly, examining interface waves in a discontinuous medium. We end the section with a demonstration of stability by considering a domain with intricate internal geometry which would be impossible to resolve using a curvilinear FD method. For time stepping, we employ a fourth order accurate modified equation approach following Gilbert and Joly [13].
@@ 4.1 Spectrum analysis
We shall expand on this by comparing the function above to another intuitive measure, the Fröbenius norm of $\Pi^{g}_{u2\tilde{u}}\Pi^{b}_{\tilde{u}2u}$ , given by
@@
as we want the spectrum to be as close to the identity matrix as possible. To optimize accuracy, we also minimize the discrete $l^{2}$-errors of leading order. This means that for an operator with a closure accuracy of $p-2$, we minimize the sum of $l^{2}$ projection errors up to order $p-1$. Initially, we minimize the spectrum, and then utilize the resulting value of the parameters as an initial step for subsequent accuracy minimization. However, it is important to note that the order in which we perform these steps is arbitrary. Therefore, it is worthwhile to explore the impact this sequencing has on the overall discretization. We remark that the optimization of free parameters in high order SBP operators was considered in [11].
To perform the minimization, we use built-in algorithms in MATLAB. We use two derivative free minimization techniques; fminunc and fminsearch to minimize the quantities (25)-(26). To investigate the effect of the projection on the overall discretization, we assemble the bulk matrix of the semidiscretization on a unit square with the DG interface at $y_{\Gamma}=1/5$, $\mu_{u,v}=1$, $\lambda_{u,v}=1$ and $\rho_{u,v}=1$ using the various projection operators obtained from the different measures and minimization techniques. The largest absolute eigenvalues of the bulk matrix, rescaled by $h^{2}$, are presented in Table 1. The first thing to note is that when the unoptimized projection operators are used, the system becomes very stiff and the CFL condition will render the method pratically unusable, motivating the need of minimization. But then it should be noted that for all different strategies used, we obtain eigenvalues with the same order of magnitude, with the largest discrepancy being when first optimizing spectrum followed by accuracy. In this case, the largest difference is a factor of $2$, which is not significant. We can therefore conclude that all choices above are suitable, but we shall resort to use the operators which resulted in the largest absolute eigenvalue of $4.431\times 10^{3}$.
@@
$\|P_{u2\tilde{u}}P_{\tilde{u}2u}\|_{E}$ $4.445\times 10^{3}$ $4.431\times 10^{3}$ $4.445\times 10^{3}$ $9.110\times 10^{3}$
$\|P_{u2\tilde{u}}P_{\tilde{u}2u}\|_{F}$ $4.438\times 10^{3}$ $4.592\times 10^{3}$ $6.040\times 10^{3}$ $4.847\times 10^{3}$
Unoptimized $1.750\times 10^{8}$
@@
The Cartesian grid in the finite difference method is constructed independently from the unstructured grid of the DG method. Thus, on the interface, a grid point from the Cartesian grid may be arbitrarily close to a node on the unstructured grid, see an illustration in Figure 2(a). In this case, we have an arbitrarily small interval on the glue grid. In the following, we investigate how this affects the projection operators and the overall discretization. More specifically, we use the same geometric setup and material parameters in the previous example. On the interface, the grid points from the Cartesian grids coincide with the nodes from the unstructured grid, except in one location where the distance is $\delta h$ where $\delta\in\{10^{-6},5\times 10^{-6},10^{-5},\cdot\cdot\cdot,5\times 10^{-2},10^{-1}\}$. We compute the largest absolute eigenvalue of the overall discretization matrix. The results are shown in Figure 2 for four different grid resolutions. From the scale of the eigenvalues, it is clear that the method is robust to non-matching FD and DG grids as the eigenvalue remains essentially constant for a fixed grid.
Figure 2: (a) An illustration of the grid on the interface (b) The largest absolute eigenvalues $\max_{i}|e_{i}|$ of the discretization matrix as a function of $\delta$. Note the small interval along the y-axis.
@@ 4.2.1 Continuous materials
We now consider a domain $\Omega$ with all material parameters chosen to be equal, $\lambda_{u}=\mu_{u}=\lambda_{v}=\mu_{v}=1$ and with unit densities $\rho_{u}=\rho_{v}=1$. Then we shall investigate the accuracy of the scheme by using the method of manufactured solution when the FD and DG interfaces match. Let $N$ denote the number of grid points on the interface. We note that even in this case, the projection operators are not identity operators because of different discrete norm in the FD and DG discretization. The solution is chosen to be given by the two fields
@@
in both the FD and DG domain. We impose traction conditions on the boundaries as induced by the chosen fields and enforce no external forcing other than that generated by the manufactured solution.
@@
$21$ $4.967\times 10^{-3}$ $5.712\times 10^{-3}$
$31$ $1.068\times 10^{-3}$ $3.790$ $6.898\times 10^{-4}$ $5.213$
$41$ $3.398\times 10^{-4}$ $3.981$ $1.600\times 10^{-4}$ $5.078$
$51$ $1.358\times 10^{-4}$ $4.109$ $5.132\times 10^{-5}$ $5.097$
$61$ $6.510\times 10^{-5}$ $4.119$ $2.045\times 10^{-5}$ $5.046$
$71$ $3.382\times 10^{-5}$ $4.147$ $9.366\times 10^{-6}$ $5.066$
@@
In our second experiment of this section, we shall instead let the overall domain $\Omega$ have its bottom boundary curved, parametrized by $f_{s}(x)=0.4x(x-1)e^{-10(x-0.5)^{2}}$ and partition it as in the previous example $\Omega_{u}=[0,1]\times[1,1/5]\times[1/5,y_{s}]$ , where $y_{s}$ denotes the function $f_{s}$ evaluated along $[0,1]$. Moreover, we use the same material parameters and densities as in the previous experiment, but let the solution be given by the fields
@@
in both FD and DG domains. Again we impose traction conditions on the boundaries as induced by the chosen fields and enforce only the forcing generated by the chosen fields. The resulting fourth and sixth order errors are presented together with the convergence orders in Table 3 and the overall behavior is illustrated in Fig. 4(b). We again note that both the fourth and sixth order methods result in an optimal convergence order even in the presence of a curved boundary, which would have required the use of more computationally expensive reference mappings if done in a pure SBP-FD setting.
@@
$21$ $4.375\times 10^{-4}$ $2.516\times 10^{-4}$
$31$ $8.229\times 10^{-5}$ $4.121$ $3.350\times 10^{-5}$ $4.973$
$41$ $2.544\times 10^{-5}$ $4.080$ $7.841\times 10^{-6}$ $5.047$
$51$ $1.074\times 10^{-5}$ $3.864$ $2.572\times 10^{-6}$ $4.994$
$61$ $5.248\times 10^{-6}$ $3.928$ $1.040\times 10^{-6}$ $4.963$
$71$ $2.754\times 10^{-6}$ $4.183$ $4.787\times 10^{-7}$ $5.037$
@@
Figure 4: (a) A coarse discretization of the domain demonstrating the linear approximation of the southern boundary given by $f_{s}(x)$ (b) The discrete error at final time $T=1$ in $||\cdot||_{h}$-norm when the material is continuous and the southern boundary is curvilinear and given by $f_{s}(x)=0.4x(x-1)e^{-10(x-0.5)^{2}}$ for both the fourth- and sixth-order methods.
@@ 4.2.2 Interface waves in discontinuous materials
If the domain $\Omega$ consists of two subdomains $\Omega_{u}$ and $\Omega_{v}$ having different material parameters, we can have waves localizing to the interface $\Gamma$ between the two. A common example of these are so-called Stoneley waves which propagate periodically in the $x$-direction with a velocity $c_{s}$ and decay exponentially from the interface. They have pressure and shear wave velocities,
$\alpha_{u}=\xi\sqrt{\frac{2\mu_{u}+\lambda_{u}}{\rho_{u}}},\quad\beta_{u}=\xi\sqrt{\frac{\mu_{u}}{\rho_{u}}},\quad\alpha_{v}=\xi\sqrt{\frac{2\mu_{v}+\lambda_{v}}{\rho_{v}}},\quad\beta_{v}=\xi\sqrt{\frac{\mu_{v}}{\rho_{v}}},$
where $\xi>0$ is a scaling factor used to fix the wavelength. Assuming that $\Omega_{u}=[0,1]\times[y_{\Gamma},1]$ and $\Omega_{v}=[0,1]\times[0,y_{\Gamma}]$ for some suitable $10<y_{\Gamma}<$ 1 we can write the waves in $\Omega_{u}$ on the following explicit form
@@
where $\gamma_{u}=\sqrt{1-c_{s}^{2}/\alpha_{u}^{2}},\eta_{u}=\sqrt{1-c_{s}^{2}/\beta_{u}^{2}}$ and in the domain $\Omega_{v}$ we instead have
@@
with $\gamma_{v}=\sqrt{1-c_{s}^{2}/\alpha_{v}^{2}}$ and $\eta_{v}=\sqrt{1-c_{s}^{2}/\beta_{v}^{2}}$. To determine the coefficients $A,B,C$ and $D$ and $c_{s}$ we substitute (27) and (28) into the interface conditions (4)-(5) obtaining a linear system
$\mathcal{K}[A,B,C,D]^{T}=[0,0,0,0]^{T}$ (29)
@@
In this experiment, we use the same computational domain as previously, $\Omega_{u}=[0,1]\times[1/5,1]$, $\Omega_{v}=[0,1]\times[0,1/5]$ and fix the wavelength to $1$ by setting $\xi=2\pi$. The nodes along the FD-DG interface are chosen to match and be given by $N$ as in the previous experiments. The material parameters used, together with the resulting amplitudes $B,C$,$D$ and wave speed $c_{s}$ are presented in Table 4.
@@
Letting the boundary traction and interior follow from the manufactured solutions (27)-(28), we obtain the errors and EOC’s are presented in Table 5 and illustrated in Figure 5(c), while the resulting fields are demonstrated at the first time-step in Figure 5(a)-5(b). While the convergence orders follow the same trend as in previous examples we note that the overall errors are larger in magnitude. This should at least partly be attributed to the large scaling used in the manufactured solution.
@@
$21$ $1.845\times 10^{-1}$- $1.849\times 10^{-1}$
$31$ $3.251\times 10^{-2}$ $4.282$ $2.433\times 10^{-2}$ $5.002$
$41$ $9.962\times 10^{-3}$ $4.111$ $5.950\times 10^{-3}$ $4.895$
$51$ $3.999\times 10^{-3}$ $4.089$ $1.942\times 10^{-3}$ $5.016$
$61$ $1.897\times 10^{-3}$ $4.090$ $7.809\times 10^{-4}$ $4.998$
$71$ $1.005\times 10^{-4}$ $4.121$ $3.617\times 10^{-4}$ $4.992$

### v2

@@ 3.2 Stability of the coupled method
It is proven in A that the conditions above are met for the choice $\chi_{2}=2\mu_{u}+\lambda_{u}$ if the penalty parameter $\beta_{v}$ satisfies the bound $\beta_{u}>\frac{1}{4C^{\prime}}$, as desired. Hence we have proven that (25) is non-negative and therefore we can conclude that $U_{2}\geq 0$ under the assumption $\beta_{u}\geq\frac{1}{4C^{\prime}}$ which together with (22) being non-negative if $\alpha_{v}\geq\frac{1}{4C}$ proves that
$\frac{1}{2}\frac{d}{dt}(\rho_{u}\boldsymbol{u}_{t}^{T}\mathcal{H}\boldsymbol{u}_{t}+(\rho_{v}\boldsymbol{v}_{t},\boldsymbol{v}_{t})_{\Omega_{v}}+U_{1}+U_{2})=0,$
@@
It should be noted that the proof will fail at (16) if we remove the assumption of full compatibility as we cannot obtain a total time derivative. This assumption can be relaxed for the wave equation [1].
@@ 4 Numerical results
In this section, we present numerical examples to demonstrate accuracy and stability properties of the FD-DG scheme developed for the elastic wave equation. [^1] The MATLAB codes used to perform the experiments are available at https://github.com/AndreasGranath/Elastic-Waves-FD-DG. First, we examine how the selection of free parameters influences the scheme’s stability and time-step restriction. Then, we conduct a series of benchmarking examples to demonstrate the stability and convergence properties of the developed method. This is demonstrated in three scenarios: first, utilizing a manufactured solution in a homogeneous medium with straight boundary and interface; then, introducing a curved boundary within the same medium; and lastly, examining interface waves in a layered medium with discontinuous material property. We end the section with a demonstration of stability by considering a domain with intricate internal geometry, which would be cumbersome to resolve using a curvilinear multiblock FD method. For time stepping, we employ a fourth order accurate modified equation approach, following Gilbert and Joly [14].
@@ 4.1 Spectrum analysis
We shall expand on this by comparing the function above to another intuitive measure, the Frobenius norm of $\Pi^{g}_{u2\tilde{u}}\Pi^{b}_{\tilde{u}2u}$ , given by
@@
as we want the spectrum to be as close to the identity matrix as possible. To optimize accuracy, we minimize the discrete $l^{2}$-errors to leading order for $\Pi^{g}_{u2\tilde{u}}$ and $\Pi^{b}_{\tilde{u}2u}$ respectively. Initially, we minimize the spectrum, and then utilize the resulting value of the parameters as an initial step for subsequent accuracy minimization. However, it is important to note that the order in which we perform these steps is arbitrary. Therefore, it is worthwhile to explore the impact this sequencing has on the overall discretization. We remark that the optimization of free parameters in high order SBP operators was considered in [12].
To perform the minimization, we use built-in algorithms in MATLAB. We use two derivative free minimization techniques; fminunc and fminsearch to minimize the quantities (26)-(27). To investigate the effect of the projection on the overall discretization, we consider the domain $[0,1]^{2}$ with traction-free boundary conditions, material properties $\mu_{u,v}=1$, $\lambda_{u,v}=1$, $\rho_{u,v}=1$ and an interface at $y_{\Gamma}=1/5$. The top block is discretized using a fourth order accurate SBP-FD method with $h=1/20$ and the bottom block is discretized by the IPDG method based on cubic basis functions. The triangulation in the bottom block is constructed such that the triangle nodes on the interface match the grid points from the top block. Then, we assemble the bulk matrix of the semidiscretization with parameters $\alpha_{u}=\beta_{u}=60$ using the various projection operators obtained from the different measures and minimization techniques. The largest absolute eigenvalues of the bulk matrix, rescaled by $h^{2}$, are presented in Table 2. The first thing to note is that when the unoptimized projection operators are used, the system becomes very stiff and the CFL condition will render the method pratically unusable, motivating the need of minimization. But then it should be noted that for all different strategies used, we obtain eigenvalues with the same order of magnitude, with the largest discrepancy being when first optimizing spectrum followed by accuracy. In this case, it is clear that the largest difference is roughly a factor of four, which is not significant. We can therefore conclude that all choices above are suitable, but we shall resort to use the operators which resulted in the smallest absolute eigenvalue of $6.104\times 10^{3}$.
@@
$\|P_{u2\tilde{u}}P_{\tilde{u}2u}\|_{E}$ $6.158\times 10^{3}$ $6.104\times 10^{3}$ $2.246\times 10^{4}$ $6.147\times 10^{3}$
$\|P_{u2\tilde{u}}P_{\tilde{u}2u}\|_{F}$ $6.119\times 10^{3}$ $6.254\times 10^{3}$ $6.281\times 10^{3}$ $9.380\times 10^{3}$
Unoptimized $1.783\times 10^{8}$
@@
The Cartesian grid in the finite difference method is constructed independently from the unstructured grid of the DG method. Thus, on the interface, a grid point from the Cartesian grid may be arbitrarily close to a node on the unstructured grid, illustrated in 2(a). In this case, we have an arbitrarily small interval on the glue grid. In the following, we investigate how this affects the projection operators and the overall discretization. More specifically, we use the same geometric setup and material parameters in the previous example. On the interface, the grid points from the Cartesian grids coincide with the nodes from the unstructured grid, except in one location where the distance is $\delta h$ where $\delta\in\{10^{-6},5\times 10^{-6},10^{-5},\cdot\cdot\cdot,5\times 10^{-2},10^{-1}\}$. We compute the largest absolute eigenvalue of the overall discretization matrix. The results are shown in Figure 2 for four different grid resolutions. From the scale of the eigenvalues, it is clear that the method is robust to non-matching FD and DG grids as the eigenvalue remains essentially constant for a fixed grid.
Figure 2: (a) An illustration of the grid on the interface (b) The largest absolute eigenvalues $\max_{i}|e_{i}|$ of the discretization matrix rescaled by $h^{2}$ as a function of $\delta$. Note the small interval along the $y$-axis.
@@ 4.2.1 Continuous materials
We again consider domain $\Omega=[0,1]^{2}$ with all material parameters chosen to be equal, $\lambda_{u}=\mu_{u}=\lambda_{v}=\mu_{v}=1$ and with unit densities $\rho_{u}=\rho_{v}=1$. Analogously, the stability parameters are chosen to be equal in both domains and given by $\alpha=\beta=60$. Then, we shall investigate the accuracy of the scheme by using the method of manufactured solution when the FD grid points match the triangle nodes of the IPDG on the interface. Let $N$ denote the number of FD grid points on the interface. We note that even in this case, the projection operators are not identity operators because of different discrete norm in the FD and DG discretization. The solution is chosen to be given by the two fields
@@
in both the FD and DG domain. We impose traction boundary conditions, and compute the boundary data and the forcing function by using the manufactured solution.
@@
$21$ $4.408\times 10^{-3}$ $5.691\times 10^{-3}$
$31$ $9.373\times 10^{-4}$ $3.818$ $6.917\times 10^{-4}$ $5.198$
$41$ $3.001\times 10^{-4}$ $3.958$ $1.597\times 10^{-4}$ $5.095$
$51$ $1.191\times 10^{-4}$ $4.139$ $5.130\times 10^{-5}$ $5.089$
$61$ $5.633\times 10^{-5}$ $4.109$ $2.042\times 10^{-5}$ $5.052$
$71$ $2.989\times 10^{-5}$ $4.104$ $9.368\times 10^{-6}$ $5.055$
@@
When using a DG scheme based on polynomial degree higher than one, there are interior nodes on the triangles. We now perform the same experiment but on a mesh where the FD grid points match all nodes on the triangle edges on the interface, see an illustration in Fig. 4(a). The discrete errors in $\|\cdot\|_{h}$-norm are presented in Table 4 and Fig. 4(b). We observe that the scheme maintains the expected convergence rate of $\min\{2p,p+2\}$ even on very fine meshes. We have not observed a clear loss of a half order in convergence rate as one would expect to happen for the SBP-FD discretization, demonstrated on a curvilinear mesh in [2].
Figure 4: (a) An illustration of the mesh structure used for mathing degrees of freedom with $p=3$, corresponding to $N=31$. (b) Numerical errors in $\|\cdot\|_{h}$ norm of the fourth and sixth order experiments for matching degrees of freedom along the interface.
$N$ Error $p=2$ EOC $p=2$ $N$ Error $p=3$ EOC $p=3$
$271$ $1.372\times 10^{-7}$- $361$ $2.578\times 10^{-9}$-
$301$ $9.020\times 10^{-8}$ $3.980$ $401$ $1.517\times 10^{-9}$ $5.031$
$331$ $5.977\times 10^{-8}$ $4.317$ $441$ $9.422\times 10^{-10}$ $5.002$
$361$ $4.223\times 10^{-8}$ $3.992$ $481$ $6.122\times 10^{-10}$ $4.881$
$391$ $3.094\times 10^{-8}$ $3.886$ $521$ $4.149\times 10^{-10}$ $4.939$
$421$ $2.279\times 10^{-8}$ $4.125$ $561$ $2.841\times 10^{-10}$ $5.110$
$451$ $1.784\times 10^{-8}$ $3.554$ $601$ $2.048\times 10^{-10}$ $4.745$
$481$ $1.396\times 10^{-8}$ $3.791$ $641$ $1.469\times 10^{-10}$ $5.147$
Least-square EOC 3.985 Least-square EOC 4.965
Table 4: Discrete errors in $||\cdot||_{h}$-norm with associated EOCs at final time $T=1$ in the case when the FD and DG degrees of freedom align. Here, $N$ denotes the number of FD grid points on the interface.
Finally, we show in Table 5 the number of degrees of freedom (DOFs) in the FD and DG schemes with $N=71$ FD grid points on the interface, which match the triangle nodes from the IPDG discretization. We also present the number of nonzero entries in the corresponding discretization matrices when $p=2$. From this, it is clear that the number of nonzeros in the discretization matrix for the DG domain is much larger than that for the FD domain, despite being a quarter of the size of the FD domain.
$N$ FD DOFs FD nonzeros DG DOFs DG nonzeros
$71$ $5041$ $247978$ $22440$ $3523200$
Table 5: Number of DOFs in the FD and DG domains together with the number of non-zero entries in the corresponding discretization matrices.
In our last experiment of this subsection, we shall instead let the overall domain $\Omega$ have its bottom boundary curved, parametrized by $f_{s}(x)=0.4x(x-1)e^{-10(x-0.5)^{2}}$ and partition it as in the previous example $\Omega_{u}=[0,1]\times[1,1/5]\times[1/5,y_{s}]$ , where $y_{s}$ denotes the function $f_{s}$ evaluated along $[0,1]$. Moreover, we use the same material parameters and densities as in the previous experiment, but let the solution be given by the fields
@@
in both FD and DG domains. Again we let the boundary traction and forcing be determined by the chosen manufactured solution. The resulting fourth and sixth order errors are presented together with the convergence orders in Table 6 and the overall behavior is illustrated in Fig. 5(b). We again note that both the fourth and sixth order methods result in an optimal convergence order even in the presence of a curved boundary.
@@
$21$ $2.936\times 10^{-4}$ $1.215\times 10^{-4}$
$31$ $5.771\times 10^{-5}$ $4.012$ $1.509\times 10^{-5}$ $5.144$
$41$ $1.823\times 10^{-5}$ $4.004$ $3.413\times 10^{-6}$ $5.168$
$51$ $7.898\times 10^{-6}$ $3.750$ $1.08\times 10^{-6}$ $5.134$
$61$ $3.796\times 10^{-6}$ $4.018$ $4.446\times 10^{-7}$ $4.895$
$71$ $2.0307\times 10^{-6}$ $4.058$ $2.014\times 10^{-7}$ $5.137$
@@
Figure 5: (a) A coarse discretization of the domain demonstrating the linear approximation of the southern boundary given by $f_{s}(x)$ (b) The discrete error at final time $T=1$ in $||\cdot||_{h}$-norm when the material is continuous and the bottom boundary of the DG domain is curvilinear, given by $f_{s}(x)=0.4x(x-1)e^{-10(x-0.5)^{2}}$ for both the fourth- and sixth-order methods.
@@ 4.2.2 Interface waves in discontinuous materials
If the domain $\Omega$ consists of two subdomains $\Omega_{u}$ and $\Omega_{v}$ having different material parameters, we can have waves localizing to the interface $\Gamma$ between the two. A common example of these are so-called Stoneley waves which propagate periodically in the $x$-direction with a velocity $c_{s}$ and decay exponentially from the interface. The components of the Stoneley waves propagate with pressure and shear wave velocities
$\alpha_{u}=\sqrt{\frac{2\mu_{u}+\lambda_{u}}{\rho_{u}}},\quad\beta_{u}=\sqrt{\frac{\mu_{u}}{\rho_{u}}},\quad\alpha_{v}=\sqrt{\frac{2\mu_{v}+\lambda_{v}}{\rho_{v}}},\quad\beta_{v}=\sqrt{\frac{\mu_{v}}{\rho_{v}}},$
and wave number $\xi>0$ Assuming that $\Omega_{u}=[0,1]\times[y_{\Gamma},1]$ and $\Omega_{v}=[0,1]\times[0,y_{\Gamma}]$ for some suitable $10<y_{\Gamma}<$ 1 we can write the waves in $\Omega_{u}$ on the following explicit form
@@
where $\gamma_{u}=\frac{\xi}{\alpha_{u}}\sqrt{\alpha_{u}^{2}-c_{s}^{2}},\eta_{u}=\frac{\xi}{\beta_{u}}\sqrt{\beta_{u}^{2}-c_{s}^{2}}$ and in the domain $\Omega_{v}$ we instead have
@@
with $\gamma_{v}=\frac{\xi}{\alpha_{v}}\sqrt{\alpha_{v}^{2}-c_{s}^{2}}$ and $\eta_{v}=\frac{\xi}{\beta_{v}}\sqrt{\beta_{v}^{2}-c_{s}^{2}}$. To determine the coefficients $A,B,C$ and $D$ and $c_{s}$ we substitute (28) and (29) into the interface conditions (4)-(5) obtaining a linear system
$\mathcal{K}[A,B,C,D]^{T}=[0,0,0,0]^{T},$ (30)
@@
In this experiment, we use the same computational domain as previously, $\Omega_{u}=[0,1]\times[1/5,1]$, $\Omega_{v}=[0,1]\times[0,1/5]$, fix the wavelength to $1$ by setting $\xi=2\pi$ and let the stability parameters be given by $\alpha=\beta=150/c_{s}$. The nodes along the FD-DG interface are chosen to match and be given by $N$ as in the previous experiments. The material parameters used, together with the resulting amplitudes $B,C$,$D$ and wave speed $c_{s}$ are presented in Table 7.
@@
Letting the boundary traction and forcingr follow from the manufactured solutions (28)-(29), we obtain the errors and convergence orders presented in Table 8, with the overall behaviour illustrated in Figure 6(c). The resulting fields at the first time-step are illustrated in Figure 6(a)-6(b). While the convergence orders follow the same trend as in previous examples we note that the overall errors are larger in magnitude. This should at least partly be attributed to the large scaling used in the manufactured solution.
@@
$81$ $5.587\times 10^{-4}$- $1.870\times 10^{-4}$
$91$ $3.397\times 10^{-4}$ $4.222$ $1.041\times 10^{-4}$ $4.968$
$101$ $2.232\times 10^{-4}$ $3.988$ $6.173\times 10^{-5}$ $4.967$
$111$ $1.538\times 10^{-4}$ $3.906$ $3.838\times 10^{-5}$ $4.986$
$121$ $1.084\times 10^{-4}$ $4.021$ $2.488\times 10^{-5}$ $4.979$
$131$ $7.846\times 10^{-5}$ $4.037$ $1.670\times 10^{-5}$ $4.978$